OpenStudy (anonymous):

Which recursive formula can be used to generate the sequence shown, where f(1) = 9.6 and n > 1? 9.6, –4.8, 2.4, –1.2, 0.6, ...

OpenStudy (retireed):

I had to look up what a recursive formula is and in simple terms it appears that a recursive formula is when previous number in the sequence is used to generate the next answer and then that answer is used to generate the next answer. initial previous answer some formula new answer a(1) a(n-1) (a(n-1) x 5) - 12 a(n) n (a(n-1)*5)-12 1 5 = a(1) 2 13 = a(2) 3 53 = a(3) 4 253 = a(4) 5 1253 = a(5) 6 6253 = a(6) 7 31253 = a(7) What is the pattern you see in your sequence of numbers?